# Near-Optimal Very Low Rank Approximation at Sublinear Cost

### v16

@@ Abstract
Eckart-Young’s theorem implies that even optimal rank-$r$ approximation of an $m\times n$ matrix $M$ for $m\geq n$ is only interesting if there is a gap in the segment of the $r+1$ largest singular values of $M$. If $M$ also has a gap in the segment of the next $p$ largest singular values, then one can compute its rank-$r$ approximation with random sampling and oversampling in nearly linear arithmetic time. For matrices $M$ having such a double gap in their spectra and for $n$ greatly exceeding $(r+p)^{2}$, our algorithm runs in sublinear time, and for a large subclass of the class of such matrices $M$ its output rank-$r$ approximation of $M$ is near-optimal under the spectral norm. We also extend the classical iterative refinement to LRA and analyze and test the resulting algorithm.
@@
low rank approximation, singular values, gaps in the spectrum, sublinear cost algorithms, iterative refinement
@@
1 LRA of a matrix: the State of the Art and our Novelties
Low rank approximation (LRA) of a matrix is a hot topic of Numerical Linear Algebra. We first briefly recall the State of the Art, referring to [11, 33, 18] for broader expositions, and then outline our LRA algorithms.
1.1 The State of the Art
1.1.1. LRA and CUR LRA. Assume that a black box matrix $M\in\mathbb{C}^{m\times n}$ is given by an oracle (black box subroutine) for computing its product with a vector rather than by its entries; then linear operations with matrices can be reduced to linear operations with vectors. [^1] The sketch of a matrix in [36, 33] is its product with a fixed set of vectors. Then we can handle matrices of immense size that come from the study of Big Data and are too large to be accessed otherwise (see further motivations in [33, Sec. 1.3]). Furthermore, quite typically the matrices representing Big Data are close to low rank matrices [35], that is, admit LRA, with which one can operate superfast, aka at sublinear cost – by using much fewer than $mn$ scalars and flops. [^2] Superfast algorithms have been studied extensively for Toeplitz, Hankel, Vandermonde, Cauchy, and other structured matrices having small displacement rank [24], and we extend this concept to other matrices.
Now assume that $m\geq n$ (otherwise we could have worked with the $n\times m$ transpose $M^{T}$) and that we are given a positive target rank $r\ll n$ or a fixed target range [^3] Here and hereafter $a\ll b$ means that $b$ greatly exceeds $a$ in context, e.g., $a=o(b)$ as $b\mapsto\infty$. for such a rank (actually, a proper choice of such targets is a challenge [11, Sec. 4.2]) and then define LRA of $M$ by a triplet $\{X,Y,Z\}$ such that
@@ 1.1 The State of the Art
The matrices $X,Y,Z$ have $mk+k\ell+\ell n\ll mn$ entries overall, and we can multiply the matrix $XYZ$ by a vector by using less than $2(mk+k\ell+\ell n)\ll mn$ flops.
If the matrices $X$ and $Z$ are made up of the columns and rows of $M$, that is, are defined by means of pre- and post-multiplication of $M$ by selected coordinate vectors, respectively, then write $X:=C$, $Z:=R$, and $Y:=U$ and call the triplet $\{X,Y,Z\}$ a CUR LRA.
1.1.2. LRA from top SVD, SVD of an LRA, and CUR from SVD. By setting to 0 all singular values of $M$ except for its $r$ top (largest) ones, obtain a matrix $M_{r}$ of rank $r$, said to be $r$-truncation of $M$. By virtue of the Eckart-Young-Mirsky theorem [7, Thm. 2.4.8], $M_{r}$ is an optimal rank-$r$ approximation444“Rank- $r$ approximation” means “rank at most $r$ approximation” throughout. of $M$ under both spectral and Frobenius norms, that is, $X=M_{r}$ minimizes these norms of $M-X$ over all matrices $X$ of rank $r$.
Computation of SVD of $M$ involves memory space for more than $mn$ scalars and order of $mn^{2}$ floating point operations (flops) (see [7, Fig. 8.6.1]). This is too expensive for computing LRA of a large size matrix, while conversely, given a triplet $\{X,Y,Z\}$ satisfying (1.1) and (1.2), we can compute SVD of $XYZ$ and then its CUR LRA by using $(m+k)k+(n+\ell)\ell$ scalars, $O(mk^{2}+n\ell^{2})$ flops and $(m+n+r)r$ scalars, $O(mr^{2})$ flops, respectively (see Appendices B and C.2, respectively).
1.1.3. LRA via random sampling and oversampling. Random sampling with oversampling enables fast computation of LRA (see [3, 17, 11, 36, 33, 20], and the references therein). For any $M\in\mathbb{C}^{m\times n}$ and target rank $k<n\leq m$, random sampling HMT algorithm of [11] successively generates a random test matrix $H\in\mathbb{C}^{n\times(k+p)}$ for some positive oversampling integer $p$, usually a small constant, computes the product $MH$, and computes rank-$k$ approximation of $M$.
@@
“In practice, the target rank $k$ is rarely known in advance. Randomized algorithms are usually implemented in anadaptive fashion where the number of samples is increased until the error norm satisfies the desired tolerance” (see [11, Sec. 4.2]).
@@
In the best studied case the test matrix $H$ is random Gaussian, filled with independent standard Gaussian (normal) random variables. In that case an expected Frobenius error norm of the output LRA is within a factor of $\epsilon=(1+\frac{k}{p-1})^{1/2}$ from the optimal value $||M-M_{r}||_{F}$ [11, Thm. 10.5], and the algorithm uses orders of $(m+n)(k+p)$ scalars and of $(k+p)mn$ flops. The parameter $p$ and the size of the test matrix $H$ explode as $\epsilon\mapsto 0$, but LRA with $\epsilon$ as large as 1/2 or 1 is still valuable.
@@
In other LRA algorithms (e.g., in [4, 3, 33, 20]) $M$ is both pre- and post-multiplied by random test matrices $F$ and $H$, respectively. In particular, the TYUC algorithm of [33], modifying those of [3, Thm. 4.7] and [3, Thm. 4.3, display 1], and the generalized Nyström approximation GN of [20] first compute the products $FM$ and $MH$ for random test matrices $F\in\mathbb{C}^{k\times m}$ and $H\in\mathbb{C}^{n\times\ell}$ for $k\leq\ell$, say, for $\ell=2k$, and then compute rank-$k$ approximation of $M$ with the expected Frobenius error norm within a factor of 2 from optimal in the case of random Gaussian test matrices $F$ and $H$ (namely, the norm does not exceed the optimal one for rank-$\frac{k}{2}$ approximation). For $\ell=2k$ the TYUC and GN algorithms use about $(m+n)\ell$ scalars (or slightly less in [20]) and $O(\ell mn)$ flops.
@@
The computation becomes a little more costly for supporting the same LRA error bounds under the spectral rather than Frobenius norm (see [11, Secs. 10 and 11], [36, Ch. 6.2], [5, Thm. 3.4]).
1.1.4. Superfast LRA for special subclasses of matrices. For some special but large and important classes of matrices, most notably for Symmetric Positive Semidefinite (SPSD) matrices but also for Distance matrices and a subclass of Cauchy matrices, the superfast randomized algorithms of [19, 1, 12, 14, 2] compute LRA whose expected Frobenius error norm is compared well with the optimal bound. The mere existence of such LRA algorithms is surprising for $n\times n$ SPSD matrices, whose class includes matrices $A^{*}A$, defined by nonsingular $n\times n$ matrices $A$ with $n^{2}$ independent entries. Notice that the superfast algorithm of [15, Part III] for SPSD matrices is deterministic, but its spectral and Frobenius output error norms exceed their optimal values by at most a factor of $n$.
Unfortunately, one cannot compute LRA superfast for every matrix: already for an input family of $mn+1$ matrices having size $m\times n$ and rank at most 1, any superfast algorithm at best halves the large error of the trivial approximation by the matrix filled with 0s (see Appendix A or, say, [26, 27, 28]). If randomization is applied, such an algorithm still at best halves the error of the trivial approximation with a probability not close to 0.
1.2 Our LRA algorithm
@@
Hereafter $\sigma_{j}(M)$ denotes the $j$th largest singular value of an $m\times n$ matrix $M$, $j=1,2,\dots$; $\sigma_{j}(M)=0$ for $j>\operatorname{rank}(M)$, and let $M$ be an $(r,\rho)$-matrix, that is, let it have gaps in both segments $[\sigma_{1}(M),\sigma_{r+1}(M)]$ and $[\sigma_{r+1}(M),\sigma_{\rho+1}(M)]$ of its singular values for $r<\rho\leq n\leq m$. We seek rank-$r$ approximation of $(r,\rho)$-matrices $M$ of size $m\times n$ such that $r<\rho\ll\sqrt{n}$. Our algorithm is superfast and, for a large subclass of $(r,\rho)$-matrices $M$, its output very low rank approximation (VLRA) is near-optimal under spectral norm.
@@
We rely on combining three surprisingly simple observations, namely, that (i) random sampling with oversampling is the only stage where HMT, TYUC, and GN algorithms are not superfast for VLRA; (ii) we can perform this stage superfast by using sparse test matrices (in particular abridged SRHT matrices of Appendix D), and (iii) while the accuracy of the VLRA output in step (ii) tends to be unsatisfactory, we can readily improve it to a near-optimal level for a large subclass of $(r,\rho)$-matrices $M$. Our LRA Alg. 2.1 of Sec. 2 relies on these observations.
@@
We compute at first (1) a crude but reasonably close rank-$\rho$ approximation $M(\rho)$ of $M$ and then (2) the rank-$r$ truncation $M(\rho)_{r}$ of $M(\rho)$, which we output as rank-$r$ approximation of $M$.
LRA via random sampling with oversampling fits this outline, but based on above observations (i)–(iii) we run the algorithm superfast for $\rho\ll\sqrt{n}$ and argue that its output VLRA is near-optimal for a large subclass of $(r,\rho)$-matrices.
We can perform stage (1) of the algorithm superfast by applying any of the HMT, TYUC and GN algorithms for $k:=\rho\ll\sqrt{n}$ with sparse test matrices such that their multiplication by $M$ is superfast (see [13], [33, Sec. 3.9], and our Sec. 3).
Apart from the stage of multiplication by test matrices, the algorithm HMT uses $(k+p)m$ scalars and $O((k+p)^{2}m)$ flops, while the TYUC and GN algorithms use about $mk+n\ell$ and $mk+n\ell+1.5kl$ scalars, respectively, and $O((m+k)k\ell)$ flops, for $k\leq\ell$, and obtain Moore-Penrose pseudoinverses of a $k\times\ell$ matrices ${\bf\Psi Q}$ in [33] and $R_{\epsilon}$ in [20] by using $O(k^{2}\ell)$ flops (see [7, Fig. 8.6.1]). These computations are superfast where $(k+p)^{2}\ll n$ and $k^{2}\ll n$, respectively, for $\ell\leq 2k$, say.
At Stage (2) we apply Alg. B.1. It computes SVD and $r$-truncation of $M(\rho)$ by using about $(m+n)\rho$ scalars and $O(mr^{2})$ flops, that is, $\ll m\sqrt{n}$ scalars and $\ll mn$ flops for $\rho\ll\sqrt{n}$.
@@
that is, Alg. 2.1 is optimal up to a factor of $1+2\frac{||M-M(\rho)||}{||M-M_{r}||}$. For $(r,\rho)$-matrices, this factor can be large only if the superfast computation of a reasonably close rank-$\rho$ approximation of $M$ at stage 1 fails. It does fail for worst case matrices $M$ (see Appendix A); moreover, empirically the output matrices of the algorithms HMT, TYUC, and GN with sparse test matrices tend to be unsatisfactory as LRA [13], [33, Sec. 3.9], but in our numerical tests in Sec. 5 accuracy of the output matrices of stage (1) of Alg. 2.1 tended to be raised to near-optimal level at stage (2).
@@ 1.3 Iterative LRA
In Sec. 4 we apply our LRA algorithm to the output error matrix, hoping to decrease its norm if it is too large. Recursively we extend to LRA the popular techniques of iterative refinement and approximate $M$ with a sum of computed LRAs. The algorithm is essentially reduced to recursive applications of Alg. 2.1, but the rank of computed sums and hence the cost of their computation grow fast as recursive steps advance. We counter this deficiency by recursively compressing the sums into rank-$r$ matrices, although this limits output accuracy of computed LRAs. The resulting iterative LRA algorithm is a natural extension to LRA of an efficient popular technique of matrix computations; we specify, analyze, and test it in Secs. 4 and 5.
We devised this algorithm to complement and strengthen Alg. 2.1, but it is essentially reduced to recursive applications of Alg. 2.1. Our upper estimates for the accuracy of both algorithms are close to one another, although both can be quite pessimistic (see Remarks 2.1 and 4.2).
@@ 1.4 Related works
Our study was prompted by extensive work on LRA in the last decades (see [11, 33, 18] and the bibliography therein) and in particular by fast LRA based on random sampling and oversampling. To accelerate these algorithms we use abridged SRHT matrices studied in the papers [26, 27, 28], which in turn extends the earlier study of Gaussian elimination without pivoting in [29, 31].
Our iterative LRA extends the algorithms of [16, 25], which in turn extend to LRA the classical technique of iterative refinement, particularly popular for the solution of a linear system of equations (see [32, Secs. 3.3.4 and 4.2.5], [10, Ch. 12 and Sec. 20.5], [7, Secs. 3.5.3 and 5.3.8]) but also widely applied in various other linear and nonlinear matrix computations (see [32, page 223 and the references on page 225]). Its extension to LRA is not straightforward because one must apply re-compression to block recursive increase of the rank of the computed approximations; in a sense we extend re-compression of [22, 30].
Actually, we first devised iterative refinement of LRA that incorporated our oversampling technique and then noticed that that technique alone can be competitive or superior.

### v17

# Iterative Refinement and Oversampling for Low Rank Approximation

@@ Abstract
Iterative refinement is particularly popular for numerical solution of linear systems of equations. We extend it to Low Rank Approximation of a matrix (LRA) and observe close link of the resulting algorithm to oversampling techniques, commonly used in randomized LRA algorithms. We elaborate upon this link and revisit oversampling and some efficient randomized LRA algorithms. Applied with sparse sketch matrices they run significantly faster and in particular yield Very Low Rank Approximation (VLRA) at sublinear cost, using much fewer scalars and flops than the input matrix has entries. This is achieved at the price of deterioration of output accuracy, but according to our formal and empirical study subsequent oversampling improves accuracy to near-optimal level under the spectral norm for a large sub-class of matrices with fast decaying spectra of singular values.
@@
low rank approximation, iterative refinement, oversampling, sparse sketches, singular values, gaps in the spectrum, sublinear cost.
@@
1 Introduction: LRA of a matrix, its iterative refinement, link to oversampling, and sparse sketches
1.1 Brief overview
Classical iterative refinement (see pointers to the bibliography in Sec. 1.6) is particularly popular for numerical solution of linear systems of equations. We naturally extend it to Low Rank Approximation of a matrix (LRA), observe its close link to LRA with oversampling, and analyze both refinement and oversampling. Here is a brief outline of our study.
Let an LRA algorithm $\mathbb{A}=\mathbb{A}_{r}$ be applied to an $m\times n$ matrix $E_{0}:=M$ and let it output a rank-$r$ approximation $Y_{0}:=X_{0}:=\mathbb{A}(E_{0})$ of $M$.  [^1] Here and hereafter “rank-$r$ approximation” means “rank at most $r$ approximation”. Apply the algorithm $\mathbb{A}$ to the error matrix $E_{1}=E_{0}-X_{0}$ and add $X_{1}:=\mathbb{A}(E_{1})$ to $X_{0}$ to obtain improved approximation $Y_{1}:=X_{0}+X_{1}$ to $M$. By repeating this step recursively $h$ times, compute the matrices $X_{i}:=\mathbb{A}(E_{i})$, $Y_{i}:=\mathbb{A}(Y_{i-1}+X_{i})$, and $E_{i+1}:=E_{i}-X_{i}$, for $i=1,2,\dots,h$. Finally output the matrix $\mathbb{A}(Y_{h})$ as a rank-$r$ approximation of $M$. The compression of higher rank approximation of a matrix into its lower rank approximation also occurs in the customary random oversampling for LRA.
We revisit random sampling LRA algorithms with oversampling, accelerate them by using sparse sketch matrices and in particular yield Very Low Rank Approximation (VLRA) at sublinear cost. We achieve this at the price of some deterioration of the accuracy of the output LRA, but according to our formal and empirical study, subsequent oversampling improves accuracy to near-optimal under the spectral norm for a large sub-class of the matrices with fast decaying spectra of singular values. So far, our study has showed no additional benefits for LRA from application of iterative refinement versus standalone oversampling.
1.2 The State of the Art of LRA
LRA is a hot topic of Numerical Linear Algebra. We first briefly recall the State of the Art, referring to [10, 27, 15] for broader expositions, and then outline our results for LRA.
1.1.1. Representation and the power of LRA. Assume that a black box matrix $M\in\mathbb{C}^{m\times n}$ is given by an oracle (black box subroutine) for computing its product with a vector rather than by its entries; then linear operations with matrices can be reduced to linear operations with vectors. [^2] The sketch of a matrix in [30, 27] is its product with a matrix of smaller size, that is, with a small set of vectors. In this way we can handle matrices of immense size that come from the study of Big Data and are too large to be accessed otherwise (see further motivations in [27, Sec. 1.3]). Furthermore, quite typically the matrices representing Big Data are close to low rank matrices [29], that is, admit LRA, with which we can operate efficiently, even where $M$ has enormous size.
Now assume that $m\geq n$ (otherwise we could have worked with the $n\times m$ transpose $M^{T}$) and that we are given a positive target rank $r<n$ or a fixed target range for such a rank, although actually a proper choice of such targets is a challenge (see Remark 1.1). Then define LRA of $M$ by a triplet $\{X,Y,Z\}$ such that
@@ 1.2 The State of the Art of LRA
1.1.2. LRA from top SVD; SVD of an LRA. By setting to 0 all singular values of $M$ except for its $r$ top (largest) ones, turn SVD of $M$ into its $r$-top SVD and turn $M$ into its $r$-truncation $M_{r}$, of rank $r$. By virtue of the Eckart-Young-Mirsky theorem [6, Thm. 2.4.8], $r$-top SVD defines an optimal rank-$r$ approximation of $M$ under both spectral and Frobenius norms, that is, $X=M_{r}$ minimizes these norms of $M-X$ over all matrices $X$ of rank $r$.
Computation of SVD of $M$ involves memory space for more than $mn$ scalars and order of $mn^{2}$ floating point operations (flops) (see [6, Fig. 8.6.1]). This is too expensive for LRA of a large size matrix, while conversely, given a triplet $\{X,Y,Z\}$ satisfying (1.1) and (1.2), we can compute SVD of $XYZ$ by using just $(m+k)k+(n+\ell)\ell$ scalars and $O(mk^{2}+n\ell^{2})$ flops (see Appendix B).
1.1.3. LRA via random sampling and oversampling. Random sampling with oversampling enables fast computation of close LRA (see [2, 14, 10, 30, 27, 16], and the references therein). For any $M\in\mathbb{C}^{m\times n}$ and target rank $k<n\leq m$, random sampling HMT algorithm of [10] successively generates a random sketch matrix $H\in\mathbb{C}^{n\times(k+p)}$ for some positive oversampling integer $p$, usually a small constant, computes the product $MH$, and computes rank-$k$ approximation of $M$.
@@
“In practice, the target rank $k$ is rarely known in advance. Randomized algorithms are usually implemented in an adaptive fashion where the number of samples is increased until the error norm satisfies the desired tolerance” (see [10, Sec. 4.2]).
@@
In the best studied case the sketch matrix $H$ is random Gaussian, filled with independent standard Gaussian (normal) random variables. In that case an expected Frobenius error norm of the output LRA is within a factor of $\epsilon=(1+\frac{k}{p-1})^{1/2}$ from the optimal value $||M-M_{r}||_{F}$ [10, Thm. 10.5], and the algorithm uses orders of $(m+n)(k+p)$ scalars and of $(k+p)mn$ flops. The integer $p$ and the size of the matrix $H$ explode as $\epsilon\mapsto 0$, but LRA for $\epsilon$ as large as 1/2 or 1 is still valuable.
@@
In other LRA algorithms (e.g., in [4, 2, 27, 16]) $M$ is both pre- and post-multiplied by random sketch matrices $F\in\mathbb{C}^{k\times m}$ and $H\in\mathbb{C}^{n\times\ell}$, respectively. In particular, the TYUC algorithm of [27], modifying those of [2, Thm. 4.7] and [2, Thm. 4.3, display 1], and the generalized Nyström approximation GN of [16] first compute the products $FM$ and $MH$ for random sketch matrices $F$ and $H$ for $k\leq\ell$, say, for $\ell=2k$, and then compute rank-$k$ approximation of $M$. In the case of random Gaussian sketch matrices $F$ and $H$ the expected Frobenius error norm of the output LRA is within a factor of 2 from optimal (namely, it does not exceed the optimal one for rank-$\frac{k}{2}$ approximation). For $\ell=2k$ the TYUC and GN algorithms use about $(m+n)\ell$ scalars (or slightly less in [16]) and $O(\ell mn)$ flops.
@@
The computation becomes a little more costly for supporting the same LRA error bounds under the spectral rather than Frobenius norm (see [10, Secs. 10 and 11], [30, Ch. 6.2], [5, Thm. 3.4]).
1.3 Acceleration of randomized LRA by using sparse sketches
@@
Hereafter $\sigma_{j}(M)$ denotes the $j$th largest singular value of an $m\times n$ matrix $M$, $j=1,2,\dots$; $\sigma_{j}(M)=0$ for $j>\operatorname{rank}(M)$, and let $M$ be an $(r,\rho)$-matrix, that is, let it have gaps in both segments $[\sigma_{1}(M),\sigma_{r+1}(M)]$ and $[\sigma_{r+1}(M),\sigma_{\rho+1}(M)]$ of its singular values for $r<\rho\leq n\leq m$. We seek rank-$r$ approximation of $(r,\rho)$-matrices $M$.
@@
We rely on combining three simple observations, namely, that (i) random sampling with oversampling is the bottleneck stage of HMT, TYUC, and GN algorithms, (ii) which we can greatly accelerate by using sparse sketch matrices (in particular abridged SRHT matrices of Appendix C), and (iii) while the accuracy of the LRA output in step (ii) tends to deteriorate with sparse sketch matrices, we readily improve it to a near-optimal level for a large subclass of $(r,\rho)$-matrices $M$.
@@
Our LRA Alg. 2.1 of Sec. 2 relies on these observations. It first computes (1) a crude rank-$\rho$ approximation $M(\rho)$ of $M$; then (2) $r$-top SVD of $M(\rho)$, output as rank-$r$ approximation of $M$.
LRA via random sampling with oversampling fits this outline.
Apart from the stage of multiplication by sketch matrices, performed at a dominated cost, the algorithm HMT uses $(k+p)m$ scalars and $O((k+p)^{2}m)$ flops, while the TYUC and GN algorithms use about $mk+n\ell$ and $mk+n\ell+1.5kl$ scalars, respectively, and $O((m+k)k\ell)$ flops, for $k\leq\ell$. Unlike the number of scalars, the bound on the number of flops decreases greatly where $k+p\ll n$ and $k\leq\ell\ll n$ for HMT and for TYUC and GN algorithms, respectively. Here and hereafter $a\ll b$ means that $b$ exceeds $a$ greatly in context, e.g., $a=o(b)$ as $b\rightarrow\infty$. In particular the algorithms HMT and TYUC/GN run at sublinear cost, that is, use much fewer than $mn$ scalars and flops, where we seek Very Low Rank Approximation (VLRA), namely, where $k+p\ll\sqrt{n}$ or $kl\ll\sqrt{n}$, respectively.
Alg. B.1 of Appendix B performs Stage (2) by using about as many scalars and $O(m\ell^{2})$ flops. We can decrease this bound to $O(mk\ell)$ (which is sublinear for $kl\ll\sqrt{n}$) by applying TUYC or GN algorithms to compute near-optimal rank-$r$ approximation of $M(\rho)$ rather than its $r$-top SVD at stage (2) of Alg. 2.1.
1.4 Accuracy of the output LRA
@@
that is, Alg. 2.1 is optimal up to a factor of $1+2\frac{||M-M(\rho)||}{||M-M_{r}||}$. For $(r,\rho)$-matrices, this factor can be large only if the computation of a reasonably close rank-$\rho$ approximation of $M$ at stage 1 fails. It does fail for worst case matrices $M$ (see Appendix A); moreover, empirically the output matrices of the algorithms HMT, TYUC, and GN with sparse sketch matrices tend to be unsatisfactory as LRA [11], [27, Sec. 3.9], but in our numerical tests in Sec. 5 accuracy of the output matrices of stage (1) of Alg. 2.1 tended to be raised at stage (2) to near-optimal level for a large sub-class of $(r,\rho)$-matrices $M$.
@@ 1.5 Iterative LRA
In Sec. 4 we apply our LRA algorithm (or we can apply any efficient LRA algorithm instead) to the output error matrix of a crude LRA computed at stage (1), to decrease its norm. Recursively we extend to LRA the popular techniques of iterative refinement and approximate $M$ with a sum of computed LRAs. We reduce the algorithm essentially to recursive applications of Alg. 2.1, but the rank of computed sums and hence the cost of their computation grow fast as recursive steps advance. We counter this deficiency by recursively compressing the sums into rank-$r$ matrices, although this limits output accuracy of computed LRAs. The resulting iterative LRA algorithm is a natural extension to LRA of an efficient popular technique of matrix computations; we specify, analyze, and test it in Secs. 4 and 5.
Our upper estimates for the accuracy of both Alg. 2.1 and iterative refinement are close to one another, although both can be quite pessimistic (see Remarks 2.1 and 4.2).
@@ 1.6 Related works
Our iterative LRA extends the algorithms of [13, 19], which in turn extend to LRA the classical technique of iterative refinement, particularly popular for the solution of a linear system of equations (see [26, Secs. 3.3.4 and 4.2.5], [9, Ch. 12 and Sec. 20.5], [6, Secs. 3.5.3 and 5.3.8]) but also widely applied in various other linear and nonlinear matrix computations (see [26, page 223 and the references on page 225]). Its extension to LRA is not straightforward because one must apply re-compression to block recursive increase of the rank of the computed approximations; in a sense we extend re-compression of [18, 24]. Such compression is commonly used for LRA with oversampling. For sparse test matrices we used abridged SRHT matrices studied in the papers [20, 21, 22], which in turn extends the earlier study of Gaussian elimination without pivoting in [23, 25].
Unlike our presentation in Secs. 1.3 and 1.5 as well as in Secs. 2 and 4, we actually first devised iterative refinement of LRA that incorporated our oversampling technique and then noticed that that technique alone can be competitive or superior.